OpenStudy (anonymous):

dy/dx = x^3(1-y) for y(0)=3; get the equation

OpenStudy (anonymous):

separation of variables gives you...\[\ln(1-y)=-{x^4 \over4}-C\]

OpenStudy (anonymous):

having trouble solving for C

OpenStudy (anonymous):

multiplied both sides by -1

OpenStudy (anonymous):

you're missing a negative sign on the left. I figured it out tho. Thx for trying

OpenStudy (ikram002p):

oops why we need to multiply both side with -1 cuz (1-y)'=-1 :) thus integral ((1-y) /-1)dy = ln (1-y) +C

OpenStudy (ikram002p):

he said he already got it besides im too sleepy xD

OpenStudy (ikram002p):

go ahead finish :P

OpenStudy (anonymous):

nope, I am not good enough to finish it, at least to the Asker, hihihi